the derivative of sec-1[1/2x2-1] with respect to root[1-x2] at x=1/2 is equal to?

Grade:

1 Answers

noogler
8 years ago
Put x=coso in sec-1[1/2x2-1] then u  get 2o put again o=cos-1x then diff with respect to x......1
similarly put x=coso in [1-x2]½  u get sino put again o=cos-1x then diff with respecttox...........2
der of 1w.r.t 2 is 2d(cos-1x)/d(sin[cos-1x])=2/(1/2)=4